if the sides of the triangle are 5.21 and 9.13 on a right triangle than what is the measurement of the other side.
Answer by Mike(39)

In this case you can simply use Pythagoras, which states that the length of both sides of a right triangle squared and then added together will give you the length of the hypotenuse squared. In mathematic terms this means: let 'a' and 'b' be the two sides of the right triangle and 'c' the hypotenuse: a^2 + b^2 = c^2. You can use this rule to determine the length of 'the other side' in your question.
5.21^2 + 9.13^2 = 110.501
110.501 = c^2
c = 10.5
Hope that helps.
Regards,
-Mike